OpenStudy (anonymous):

This simultaneous equation is stressing me out, anyone can help what I am doing wrong?! 2x + y = 19 7x + y = 29 Multiplied the 2 x7 and the 7 X2 14x + 14y = 133 14x + 2y = 58 Then take it away: 12y = 75 y = 75/12 = 6.25 Then find the X 2x + 6.25 = 19 2x = 12.75 / 2 = 6.375 When I try to work out the 2 x 6.375 + 6.25 = 19 is fine but then 7 x 6.375 + 6.25 doesn't give me 29!!! GRR

OpenStudy (anonymous):

Where you have mulltiplied the first equation by 7, you should actually have 14x+7y=133

OpenStudy (anonymous):

2x + y = 19 7x + y = 29 14x + 7y = 133

OpenStudy (anonymous):

yep safiah

OpenStudy (anonymous):

@Safia what a silly mistake, I am tired I should leave Maths for a while. Silly mistake thanks for noticing!

OpenStudy (anonymous):

no problem.Here to help :-)

OpenStudy (anonymous):

serrico - i know what u mean - i just made a mistake in my last reply - i'm too tired to carry on !!